OpenStudy (anonymous):

Convert each radian measure to degrees. Give answers to the nearest 10 minutes or tenth of a degree. 1.6

OpenStudy (anonymous):

\[\frac{1.6 }{\pi }*180=91.7{}^{\circ} \]

OpenStudy (anonymous):

Do you know why the other option is 91 degrees 40'?

OpenStudy (anonymous):

The answer units option was minutes or 1/10 degrees. \[91.7{}^{\circ}=91{}^{\circ} 42 \text{ minutes} = (91{}^{\circ} 42' <- \text{correct} \text{ notation}?) \]

OpenStudy (anonymous):

Overlooked the "nearest 10 minutes" requirement.\[91{}^{\circ} 40\text{ }\text{minutes} \]